OpenStudy (anonymous):

A sine function can be used to model light waves. Green light has a wavelength, or period, of about 520 nanometers(nm). Which equation best models green light?

OpenStudy (anonymous):

the period is always \(\frac{2\pi}{b}\)

OpenStudy (anonymous):

A. y= sin 260/pi theta B. y= sin pi/520 theta C. y= sin 520/pi theta D. y= sin pi/260 theta

OpenStudy (anonymous):

set \[\frac{2\pi}{b}=520\] solve for \(b\) in one or two steps

OpenStudy (anonymous):

Is it A or D?

OpenStudy (anonymous):

b=260 I think it's D?

OpenStudy (anonymous):

\[b=\frac{2\pi}{520}=\frac{\pi}{260}\]

OpenStudy (anonymous):

so yeah D

OpenStudy (anonymous):

Thanks so much for your help!

OpenStudy (anonymous):

yw